Can someone please help me with this one;
1019.03 = 25.57802 + 1026/(1 + x/2)^2

Answer by RAY100(1637)

1019.03 = 25.57802 + 1026/ (1+x^2)^2
.
1019.03 - 25.57802 = 1026 / (1+x^2)^2
.
993.4520 =1026/(1+x^2)^2
.
cross multiple
.
(1+x^2)^2 = 1026 / 993.4520 = 1.03276255
.
take sqrt both sides
.
(1+x^2) = 1.016
.
x^2 = .016
,
take sqrt both sides
.
x= .127472569
.
check,,,(1+.1274^2)^2 = 1.0328
.
1019.03=25.57802 + 1026/1.0328
.
1019.03 = 1019.03,,,,,ok